# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

### v2

@@ Abstract
We investigate Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), accelerometry, and temperature, and finger photoplethysmography (PPG) and temperature.
@@ Methods
We obtained wearable sensor recordings from 360 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab. Each PSG recording was manually scored and these annotations served as ground truth labels for training and evaluation of our models. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained Mamba-based models with convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings. Ensembling of model variants with similar architectures was performed.
@@ Results
Our best approach, after ensembling, attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, NREM stage 1/2 [N1/N2], NREM stage 3 [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
@@ Conclusions
Deep learning models can infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be successfully applied to data from adults attending a tertiary care sleep clinic.
@@
3Vector Institute Toronto, ON, Canada
@@
[…]
@@ 1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in evaluating individuals with and at risk for neurodegenerative diseases [1]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment. This is followed by sleep staging, the manual categorization of time periods into different sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep. EEG is the principal signal used for sleep staging [2, 3, 4].
@@ 1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may discomfort patients, resulting in sleep that is unrepresentative of that in the home environment. Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in rural or remote areas without PSG services. Furthermore, monitoring patients throughout the night and visually labelling sleep epochs is a time-consuming process that limits the throughput of sleep analyses.
In response to these challenges, wearable devices have been developed that record a subset of conventional PSG signals. Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs. On the other hand, EEG-less devices are less obtrusive and easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [7], transformers [8], and graph neural networks [9]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
@@ 1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that takes in a sequence of vectors (e.g. a time series of pulse waves, or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
Mamba [14] provides a solution to both problems as a type of state space model (SSM) [15] that combines the excellent long-range dependency modeling of SSMs with an additional time-dependent selection mechanism that allows for focusing on or ignoring particular features, as well as a hardware-friendly algorithm that provides computational efficiency. Empirically, Mamba has achieved state-of-the-art performance on supervised and unsupervised metrics using datasets of text, audio waveforms, and DNA sequences [14], while Mamba-based models have been adopted to perform automatic sleep staging using EEG and other PSG signals to high degrees of accuracy [16, 17].
@@ 1.4 Contribution
This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and temperature. The raw signals of a typical recording are plotted in Figure 1B alongside its hypnogram.
@@
A prior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance. We aim to develop a superior population model to achieve accurate sleep staging from ANNE One signals recorded from previously unseen patients.
@@
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
@@
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (41% were over the age of 65 and 46% had sleep apnea; Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. This is particularly important as age, sleep, and neurological conditions can all plausibly alter the relationship between sleep stage and autonomic, cardiac, and pulmonary function, meaning that models trained on younger adults may generalize poorly to older adults. Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [16, 17] are also noticed in this context.
@@ 2.1 Dataset
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [2]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, as were time periods prior to the sensor placement or after the sensors were taken off, leaving a total of 357 patients with 2914 hours of recording. The characteristics of the study’s participants are outlined in Figure 1C. Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Appendix A).
The machine learning task is defined from the dataset as a form of multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
@@ 2.2 Preprocessing
The ECG data were recorded at 512 Hz and preprocessed as follows. A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. For each peak, morphological features (e.g. amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters. Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes. QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
The PPG data were recorded at 128 Hz and preprocessed as follows. First, PPG signals were band-pass filtered between and downsampled to 100 Hz. Trains of physiological pulse waves were identified on the basis of relatively stable HR and amplitude. These high-quality pulse waves were normalized to a standard amplitude and duration, and clustered using HDBSCAN. For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster; beat-to-beat template matching was performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave. For each matching pulse wave, the timing of the foot and peak was noted, from which a measure of heart rate was obtained.
The pulse arrival time (PAT) is the delay from the R-peak of the ECG to the arrival of the pulse wave in the PPG. To compute PAT from the ANNE data, the time-series of valid ECG peaks and PPG feet were matched to each other by empirically considering all possible lags within the range msec and selecting that which successfully matched the greatest number of ECG peaks and PPG feet. Afterwards, for each ECG R-peak and its corresponding PPG foot, the PAT of that heartbeat was computed, generating a beat-to-beat time series of PAT.
The chest XYZ accelerometry data were recorded at 210 Hz and low-pass filtered, followed by the computation of the euclidean norm minus one (ENMO), theta (body pitch), phi (body roll) and the z-angle (the deviation from vertical of a vector perpendicular to the face of the sensor).
@@ 2.3 Model
Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3. The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs. The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs. 70K). Since Mamba is designed for long-context pattern recognition, it could potentially interact with signals more effectively than the CRNN if given them directly without windowing.
The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows (mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria), the window size (i.e. 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features. The model comprises:
@@
A convolutional component, consisting of 1D residual blocks that sequentially downsize the time-series and frequency datas’ window size axes and flatten them into an embedding vector per window. The time-series embedding, frequency embedding and scalar data are then concatenated across the feature axis to form an embedding vector per window.
@@
A recurrent component, consisting of a bidirectional three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class. The bidirectional Mamba block (Appendix B) is the concatenation of two Mamba blocks that process epochs in opposite directions, allowing for a widened receptive field to detect patterns in signals more effectively.
The RNN architecture (Figure 2B) processes a single 2D input channel with axes for the number of -second windows and number of features, where is an arbitrary value dividing 30. It comprises:
1)
An element-wise MLP with hidden layer dimensions that downsizes the data to dimensions, where the choices of and are arbitrary.
2)
A bidirectional three-layer Mamba block that processes windows sequentially.
@@
An element-wise MLP that downsizes the data into logits, which, during inference-time, can be grouped and averaged to form logits per 30-second window to comply with AASM standards.
Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches. The engineered scalar features, which have also shown discriminatory value in other previous models [30], distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix D). Thus, empirical selections of these features (Table 1) were introduced along with the time-series signals for both the CRNN and RNN.
A series of transformations were applied to further adjust the preprocessed data to the individual architectures. For the CRNN, all signals were resampled to 25 Hz and grouped to 30-second windows, resulting in data points per window. Then, depending on feature type (see Table 1),
For time-series features, the signal was left alone.
For frequency features, the log-magnitude spectrum was taken of the detrended signal over the 30-second axis.
For scalar features, the signal was reduced over the 30-second axis. Metadata, in the form of patient age and sex (male = 0, female = 1), was incorporated into the scalar features by repeating the constants times throughout the recording.
For the RNN, the accelerometry, ECG, and PPG signals were detrended, with the accelerometry signals further band-pass filtered over Hz. Log-magnitude spectrograms were taken with a window size of 100 points and no overlap, then finally concatenated with limb/chest temperature, which was resampled to 1 Hz, and sleep stage label, of which each point was repeated 30 times to 1 Hz.
RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO)1 Z* Entropy, PPG Chest Temperature Z-Angle2 Z-Angle* Entropy, PPG* Limb Temperature Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle

### v4

# Mamba-based Deep Learning Approach for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

@@ Abstract
We investigate a Mamba-based deep learning approach for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and chest temperature, and finger photoplethysmography and finger temperature.
@@ Methods
We obtained wearable sensor recordings from 357 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab. Each PSG recording was manually scored and these annotations served as ground truth labels for training and evaluation of our models. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained a Mamba-based recurrent neural network architecture on these recordings. Ensembling of model variants with similar architectures was performed.
@@ Results
After ensembling, the model attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, light NREM [N1/N2], deep NREM [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
@@ Conclusions
Our Mamba-based deep learning model can successfully infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be applied to data from adults attending a tertiary care sleep clinic.
@@
3Vector Institute for Artificial Intelligence Toronto, ON, Canada
@@
# {andrewhz.zhang, alex.hemo, r.yin}@mail.utoronto.ca
@@
Statement of Significance
Sleep is traditionally measured by in-lab polysomnography followed by visual identification of sleep stages by specialized technologists. However, the need for specialized personnel and equipment makes polysomnography difficult to scale, and poses a barrier to many older adults. To address this, we developed and evaluated in 357 patients attending a tertiary care sleep clinic a Mamba-based AI model for inferring sleep stage from the ANNE One sensor system – a pair of wearable minimally intrusive sensors placed on the chest and finger. We show good performance at distinguishing sleep stages across a range of ages and co-morbid sleep pathologies, suggesting that this may represent a scalable, accurate, minimally intrusive approach to the ambulatory assessment of sleep in clinical populations.
Graphical Abstract
![[Uncaptioned image]](figures/graphical_abstract.png)
@@ 1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders, playing an important role in evaluating individuals with and at risk for neurodegenerative diseases [iranzo2015sleep]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously by sensors attached to a sleeping patient in a laboratory environment. This is followed by sleep staging, the categorization of time periods into sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [aasm] conventions, recordings are divided into 30-second windows, or "epochs". Electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) are used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep. EEG is the principal signal used for sleep staging [aasm, rechtschaffen1968brain, imtiaz2021systematic].
@@ 1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may also discomfort patients, resulting in sleep that is unrepresentative of that in the home environment. Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in remote areas without PSG services. Furthermore, monitoring patients throughout the night and manually labelling sleep epochs is a time-inefficient process that limits the volume of sleep analyses.
In response to these challenges, wearable devices have been developed that record a subset of conventional PSG signals. Devices that record EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs. On the other hand, EEG-less devices are less obtrusive and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-laboratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [tsinalis2016automatic], recurrent neural networks (RNNs) [malafeev2018automatic], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [seo2020intra], transformers [phan2022sleeptransformer], and graph neural networks [jia2020graphsleepnet]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [fiorillo2019automated, phan2022automatic].
@@ 1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that takes in a sequence of vectors (e.g. a time series of pulse waves or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short-term memory (LSTM) [hochreiter1997long] and gated recurrent units (GRU) [cho2014learning], which are known to forget very long-range dependencies and whose architecture cannot exploit GPU acceleration, limiting the training speed. Transformers [vaswani2017attention] address these problems but scale poorly in speed for longer sequences both for training and inference.
Mamba [mamba] as a new generation of RNN provides a solution to these issues as a type of state space model (SSM) [gu2021combining] that combines the excellent long-range dependency modeling of SSMs with a time-dependent selection mechanism that allows for focusing on particular features, as well as a hardware-friendly algorithm that is computationally efficient (see S.1 for details). Empirically, Mamba has achieved state-of-the-art performance rivaling that of transformers on supervised and unsupervised metrics from datasets of text, audio waveforms, and DNA sequences [mamba], while Mamba-based models have been adopted to perform sleep staging using EEG and other PSG signals to high degrees of accuracy [zhang2024mssc, zhou2024bit].
![[Uncaptioned image]](figures/anne_data.drawio.png)
@@ 1.4 Contribution
This paper applies a Mamba-based deep learning approach to the problem of sleep staging using data from the ANNE One [sibel] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health. It is composed of a finger module, which measures photoplethysmography (PPG) and limb temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and chest temperature. The raw signals of a typical recording are plotted in Figure 1B alongside a hypnogram, a technician-labelled plot of sleep stage over time.
@@
A prior study centered on ANNE One [anne_personal] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient; in contrast, the population model, trained and evaluated on data from different patients, showed much poorer performance. We aim to develop a superior population model to achieve accurate sleep staging from ANNE One signals recorded from previously unseen patients.
@@
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [radha2019sleep, sun, sleepppg, pini, topalidis2023pulses, jones2024expert, kazemi2024improved], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss. Consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [silva, zhang2018sleep]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
@@
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (41% were over the age of 65 and 46% had sleep apnea; Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. This is particularly important as age, sleep, and neurological conditions can all plausibly alter the relationship between sleep stage and autonomic, cardiac, and pulmonary function, meaning that models trained on younger adults may generalize poorly to older adults. Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [zhang2024mssc, zhou2024bit] are also noticed in this context.
@@ 2.1 Dataset
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [aasm]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, as were time periods prior to the sensor placement or after the sensors were taken off, leaving a total of 357 patients with 2914 hours of recording. The characteristics of the study’s participants are outlined in Figure 1C. Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Figure S.2)
The machine learning task is defined as multiclass 1D-signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time to form a hypnogram as demonstrated in Figure 1B.
@@ 2.2 Preprocessing
The ECG data were recorded at 512 Hz and processed by a multilevel 1D Coiflet transform with 5 moments. Peaks in the resulting signal were identified and used to define windows; within each window, the original ECG signal was searched for local maxima and minima to determine the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. We then computed an ECG signal quality index (SQI) as follows. Trains of 5 or more physiological QRS complexes were identified on the basis of relatively stable HR (differing by no more than 10 beats per minute between consecutive beats) and amplitude (standard deviation of amplitude ). All identified ECG peaks in the recording were then clustered using HDBSCAN [mcinnes2017hdbscan] based on amplitude and morphology; peaks that clustered with the template QRS complexes were considered to have passed quality control (QC). Segments of ECG of 2 seconds on either side of these peaks were assigned an SQI of 1 while all other segments were given an SQI of 0.
The PPG data were recorded at 128 Hz, then band-pass filtered between and downsampled to 100 Hz. Trains of 5 or more physiological pulse waves were identified on the basis of relatively stable HR (differing by no more than 10 beats per minute between consecutive beats) and amplitude (differing by no more than 2-fold between beats). These high-quality pulse waves were normalized to a standard amplitude and duration, then clustered using HDBSCAN. For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster. Beat-to-beat template matching was then performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave. Pulse waves with minimum sum squared were considered to have passed QC. Segments of PPG signal were set to 1 if they were within 2 seconds of a pulse wave passing QC, and 0 otherwise.
The chest XYZ accelerometry data were recorded at 210 Hz, downsampled to 100 Hz similar to ECG, and band-pass filtered over Hz with a second-order Butterworth filter.
Log-magnitude spectrograms were taken of each signal by means of a short-time Fourier transform (STFT) with a Hamming window of size equal to the signals’ sampling rate (i.e. 100) and with no overlap, creating samples at Hz (an arbitrary value that was determined experimentally). Only the first 12 frequency bins of the XYZ data were kept as the other higher-frequency bins were made redundant by the band-pass filter. Finally, the frequency bins were concatenated with limb/chest temperature, which was recorded at 1 Hz, and sleep stage label, which was created at Hz (as labels are for 30-second windows) and upsampled to 1 Hz by repeating each label 30 times.
@@ 2.3 Model
A Mamba-based RNN model (summarized in Figure 2B) was designed to process batches of 2D tensors with axes for the number of -second windows and number of features, which include chest temperature, limb temperature, and frequency bins for ECG, PPG, and XYZ. Logits for each sleep stage are returned at each time step. A model with output classes has 70K parameters which constitute:
@@
An element-wise multi-layer perceptron (MLP) with hidden dimensions that diminishes the feature-size axis to dimensions, where and are experimentally determined.
@@
A bidirectional three-layer Mamba block. This consists of two Mamba blocks [mambapy] that process time windows sequentially in opposite directions, whose outputs are concatenated in the feature-length axis to allow for a widened receptive field that detects patterns in signals more effectively.
@@
An element-wise MLP that downsizes the feature size dimension into logits.
In accordance with AASM practice, we form 30-second window predictions by grouping 1-second logits into contiguous buckets of 30, taking the mean logit for each class, and outputting the sleep stage with the largest corresponding logit.